# Minimal criteria for continuous-variable genuine multipartite entanglement

### v1

@@ C.1 N=3
Next, for simplicity we round all elements of the latter CM to two decimal places thereby getting the matrix $\bar{\gamma}_{3}$, which slightly violates the Heisenberg uncertainty principle $\gamma+i\Omega_{3}\geq 0$ [41], Eq. (8) with $N=3$ of the main text. Therefore, we have added to the CM a small amount of thermal noise thereby creating a physical CM $𝟙\gamma_{3}=\bar{\gamma}_{3}+0.01\times\openone$ given explicitly in Eq. (78). Now, if we optimize numerically the difference (53) for the sum criterion (V.1) over the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, we arrive at $D_{S}=-0.15$ and the optimal parameters are listed in Tab. 7.
Table 7: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference (53) for the sum criterion (V.1) and the CM $\gamma_{3}$, Eq. (78).
@@
The difference (53) in the first row of Tab. 3 of the main text has been calculated for the unrounded CM $\gamma_{3}^{\prime}$, Eq. (102). If we optimize numerically the difference (53) for the sum criterion (V.1) over the parameters $\ell^{\alpha}_{ij}\in[-10,10]$, we arrive at $D_{S}=-17.87$ and the optimal parameters are listed in Tab. 8.
@@ C.2.1 Linear tree 4a
$Z_{4a}^{x}=\left(\begin{array}[]{cccc}1&-0.5&0&0\\ -0.5&0.5&-0.5&0\\ 0&-0.5&1.25&-0.5\\ 0&0&-0.5&1.0001\\ \end{array}\right)\$ (105)
@@
$Z_{4a}^{p}=\left(\begin{array}[]{cccc}1&0.5&0&0\\ 0.5&0.5&0.5&0\\ 0&0.5&1.25&0.5\\ 0&0&0.5&1.0001\\ \end{array}\right).\$ (106)
@@
$\gamma_{4a}=S_{4a}^{T}S_{4a}=\left(\begin{array}[]{cccc}1.89322&1.79444&0.87087&0.350295\\ 1.79444&2.61924&1.70931&0.530383\\ 0.87087&1.70931&1.96209&0.828527\\ 0.350295&0.530383&0.828527&1.23606\\ \end{array}\right)\oplus\left(\begin{array}[]{cccc}1.89322&-1.79444&0.87087&-0.350295\\ -1.79444&2.61924&-1.70931&0.530383\\ 0.87087&-1.70931&1.96209&-0.828527\\ -0.350295&0.530383&-0.828527&1.23606\\ \end{array}\right).$ (128)
@@ C.2.2 Star tree 4b
$Z_{4b}^{x}=\left(\begin{array}[]{cccc}4&0&0&-2\\ 0&1&0&-1\\ 0&0&4&-2\\ -2&-1&-2&3.01\\ \end{array}\right)\$ (131)
@@
$Z_{4b}^{p}=\left(\begin{array}[]{cccc}100&0&0&10\\ 0&25&0&5\\ 0&0&25&5\\ 10&5&5&3.01\\ \end{array}\right).\$ (132)
@@
$\gamma_{4b}=S_{4b}^{T}S_{4b}=\left(\begin{array}[]{cccc}5.76616&1.52878&0.765563&1.53348\\ 1.52878&8.05052&1.5276&3.0599\\ 0.765563&1.5276&3.26498&1.5323\\ 1.53348&3.0599&1.5323&2.06932\\ \end{array}\right)\oplus\left(\begin{array}[]{cccc}0.230645&0.061151&0.0612449&-0.306695\\ 0.061151&0.322021&0.122208&-0.611981\\ 0.0612449&0.122208&0.522397&-0.612921\\ -0.306695&-0.611981&-0.612921&2.06932\\ \end{array}\right).$ (166)
@@ C.3.1 Star tree 5b
$Z_{5b}^{x}=\left(\begin{array}[]{ccccc}1&0&0&0&-0.1\\ 0&0.25&0&0&-0.05\\ 0&0&1&0&-0.1\\ 0&0&0&0.25&-0.05\\ -0.1&-0.05&-0.1&-0.05&0.0401\\ \end{array}\right)\$ (169)
@@ C.3.2 Tree 5c
$Z_{5c}^{x}=\left(\begin{array}[]{ccccc}1&-0.1&0&0&0\\ -0.1&0.26&0&0&-0.05\\ 0&0&1&0&-0.1\\ 0&0&0&0.25&-0.05\\ 0&-0.05&-0.1&-0.05&0.0301\\ \end{array}\right)\$ (185)
@@
$Z_{5c}^{p}=\left(\begin{array}[]{ccccc}1&0.1&0&0&0\\ 0.1&0.26&0&0&0.05\\ 0&0&1&0&0.1\\ 0&0&0&0.25&0.05\\ 0&0.05&0.1&0.05&0.0301\\ \end{array}\right).\$ (186)
@@ C.4.2 Tree 6d
$Z_{6d}^{x}=\left(\begin{array}[]{cccccc}1&0&0&-0.1&0&0\\ 0&1&0&-0.1&0&0\\ 0&0&1&0&-0.1&0\\ -0.1&-0.1&0&0.27&0&-0.05\\ 0&0&-0.1&0&0.02&-0.01\\ 0&0&0&-0.05&-0.01&0.0201\\ \end{array}\right)\$ (219)
@@
$Z_{6d}^{p}=\left(\begin{array}[]{cccccc}1&0&0&0.1&0&0\\ 0&0.25&0&0.05&0&0\\ 0&0&1&0&0.1&0\\ 0.1&0.05&0&0.27&0&0.05\\ 0&0&0.1&0&0.02&0.01\\ 0&0&0&0.05&0.01&0.0201\\ \end{array}\right).\$ (220)
@@ C.4.3 Tree 6e
$Z_{6e}^{p}=\left(\begin{array}[]{cccccc}4&0&0&0&0.2&0\\ 0&1&0&0&0.1&0\\ 0&0&1&0&0&0.1\\ 0&0&0&1&0&0.1\\ 0.2&0.1&0&0&0.03&0.01\\ 0&0&0.1&0.1&0.01&0.0301\\ \end{array}\right).\$ (238)

### v3

@@ D.1 N=3
Next, for simplicity we round all elements of the latter CM to two decimal places thereby getting the matrix $\bar{\gamma}_{3}$, which slightly violates the Heisenberg uncertainty principle $\gamma+i\Omega_{3}\geq 0$ [54], Eq. (6) with $N=3$ of the main text. Therefore, we have added to the CM a small amount of thermal noise thereby creating a physical CM $\gamma_{3}=\bar{\gamma}_{3}+0.01\times\mathbbm{1}$ given explicitly in Eq. (78). Now, if we optimize numerically the difference (49) for the sum criterion (5.1) over the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, we arrive at $D_{S}=-0.15$ and the optimal parameters are listed in Tab. 7.
Table 7: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference (49) for the sum criterion (5.1) and the CM $\gamma_{3}$, Eq. (78).
@@
The difference (49) in the first row of Tab. 3 of the main text has been calculated for the unrounded CM $\gamma_{3}^{\prime}$, Eq. (94). If we optimize numerically the difference (49) for the sum criterion (5.1) over the parameters $\ell^{\alpha}_{ij}\in[-10,10]$, we arrive at $D_{S}=-17.87$ and the optimal parameters are listed in Tab. 8.
@@ D.2.1 Linear tree 4a
$Z_{4a}^{x}=\left(\begin{array}[]{cccc}1&-0.5&0&0\\ -0.5&0.5&-0.5&0\\ 0&-0.5&1.25&-0.5\\ 0&0&-0.5&1.0001\\ \end{array}\right)\\$
@@
$Z_{4a}^{p}=\left(\begin{array}[]{cccc}1&0.5&0&0\\ 0.5&0.5&0.5&0\\ 0&0.5&1.25&0.5\\ 0&0&0.5&1.0001\\ \end{array}\right).\\$
@@
$\gamma_{4a}=S_{4a}^{T}S_{4a}=$ $\left(\begin{array}[]{cccc}1.89322&1.79444&0.87087&0.350295\\ 1.79444&2.61924&1.70931&0.530383\\ 0.87087&1.70931&1.96209&0.828527\\ 0.350295&0.530383&0.828527&1.23606\\ \end{array}\right)$ (103)
$\oplus\left(\begin{array}[]{cccc}1.89322&-1.79444&0.87087&-0.350295\\ -1.79444&2.61924&-1.70931&0.530383\\ 0.87087&-1.70931&1.96209&-0.828527\\ -0.350295&0.530383&-0.828527&1.23606\\ \end{array}\right).$ (108)
@@ D.2.2 Star tree 4b
$Z_{4b}^{x}=\left(\begin{array}[]{cccc}4&0&0&-2\\ 0&1&0&-1\\ 0&0&4&-2\\ -2&-1&-2&3.01\\ \end{array}\right)\\$
@@
$Z_{4b}^{p}=\left(\begin{array}[]{cccc}100&0&0&10\\ 0&25&0&5\\ 0&0&25&5\\ 10&5&5&3.01\\ \end{array}\right).\\$
@@
$\gamma_{4b}=S_{4b}^{T}S_{4b}=$ $\left(\begin{array}[]{cccc}5.76616&1.52878&0.765563&1.53348\\ 1.52878&8.05052&1.5276&3.0599\\ 0.765563&1.5276&3.26498&1.5323\\ 1.53348&3.0599&1.5323&2.06932\\ \end{array}\right)$ (119)
$\oplus\left(\begin{array}[]{cccc}0.230645&0.061151&0.0612449&-0.306695\\ 0.061151&0.322021&0.122208&-0.611981\\ 0.0612449&0.122208&0.522397&-0.612921\\ -0.306695&-0.611981&-0.612921&2.06932\\ \end{array}\right).$ (124)
@@ D.3.1 Star tree 5b
$Z_{5b}^{x}=\left(\begin{array}[]{ccccc}1&0&0&0&-0.1\\ 0&0.25&0&0&-0.05\\ 0&0&1&0&-0.1\\ 0&0&0&0.25&-0.05\\ -0.1&-0.05&-0.1&-0.05&0.0401\\ \end{array}\right)$
@@ D.3.2 Tree 5c
$Z_{5c}^{x}=\left(\begin{array}[]{ccccc}1&-0.1&0&0&0\\ -0.1&0.26&0&0&-0.05\\ 0&0&1&0&-0.1\\ 0&0&0&0.25&-0.05\\ 0&-0.05&-0.1&-0.05&0.0301\\ \end{array}\right)\\$
@@
$Z_{5c}^{p}=\left(\begin{array}[]{ccccc}1&0.1&0&0&0\\ 0.1&0.26&0&0&0.05\\ 0&0&1&0&0.1\\ 0&0&0&0.25&0.05\\ 0&0.05&0.1&0.05&0.0301\\ \end{array}\right).\\$
@@ D.4.2 Tree 6d
$Z_{6d}^{x}=\left(\begin{array}[]{cccccc}1&0&0&-0.1&0&0\\ 0&1&0&-0.1&0&0\\ 0&0&1&0&-0.1&0\\ -0.1&-0.1&0&0.27&0&-0.05\\ 0&0&-0.1&0&0.02&-0.01\\ 0&0&0&-0.05&-0.01&0.0201\\ \end{array}\right)$
@@
$Z_{6d}^{p}=\left(\begin{array}[]{cccccc}1&0&0&0.1&0&0\\ 0&0.25&0&0.05&0&0\\ 0&0&1&0&0.1&0\\ 0.1&0.05&0&0.27&0&0.05\\ 0&0&0.1&0&0.02&0.01\\ 0&0&0&0.05&0.01&0.0201\\ \end{array}\right).$
@@ D.4.3 Tree 6e
$Z_{6e}^{p}=\left(\begin{array}[]{cccccc}4&0&0&0&0.2&0\\ 0&1&0&0&0.1&0\\ 0&0&1&0&0&0.1\\ 0&0&0&1&0&0.1\\ 0.2&0.1&0&0&0.03&0.01\\ 0&0&0.1&0.1&0.01&0.0301\\ \end{array}\right).\\$
@@ D.4.4 Tree 6f
Appendix E Numerical example of robust GME state
This section contains CM of state found by Gaussian analog [43] of the iterative algorithm [41], as is done in Sec. 5.1.2. In contrast to the states with CMs $\gamma_{1}$ and $\gamma_{2}$ presented there, we used additional constraint restricting the amount of squeezing needed to prepare the state with given CM to maximally $5\;\mathrm{dB}$. The state with CM
$\gamma_{7}=$ $\left(\begin{array}[]{ccc}3.96432&3.33486&1.66686\\ 3.33486&3.60552&1.84641\\ 1.66686&1.84641&1.27596\end{array}\right)$ (152)
$\oplus\left(\begin{array}[]{ccc}3.48332&-2.61754&3.45572\\ -2.61754&3.13354&-3.379668\\ 3.45572&-3.79668&7.73756\end{array}\right)$ (156)
is detected as GME by criterion (5.1) with the difference $D_{P}=-0.111$ over the parameters $\ell^{\alpha}_{ij}\in[-1,1]$ given in Tab. 11.
Table 11: Rounded values of the parameters $\ell^{\alpha}_{ij}\in[-1,1]$, $\alpha=x,p$, used to calculate the difference (48) for the product criterion (5.1) and the CM $\gamma_{7}$, Eq. (152).
$\ell^{x}_{11}$ 0.67 $\ell^{p}_{11}$ 0.82
$\ell^{x}_{12}$ -0.76 $\ell^{p}_{12}$ 1.00
$\ell^{x}_{22}$ 0.68 $\ell^{p}_{22}$ 1.00
$\ell^{x}_{23}$ -1.00 $\ell^{p}_{23}$ 0.61
$\ell^{x}_{33}$ 0.01 $\ell^{p}_{33}$ 0.01